OpenStudy (anonymous):

The radius of a hydrogen atom is about 0.37Å, and the average radius of the earth's orbit around the sun is about 1.5x10^8 km. Find the ratio of the average radius of the earth's orbit to the radius of the Hydrogen atom.

OpenStudy (anonymous):

To find the ratio you have to divide the two radius with each other. Make sure they both have the same unit.

OpenStudy (anonymous):

1:2.46x10^-22 I believe

OpenStudy (anonymous):

Hint: An Angstrom is equal to \(\bf 1.0 \times 10^{-10} \ m \)

OpenStudy (anonymous):

\[\large \bf \frac{ 0.37 \times 10^{-10} \ m}{ 1.5 \times 10^8 \ m } \approx 2.5 \times 10^{-19}\]